OpenStudy (anonymous):

f(x)= 4x^2+10x+10 Express f(x) in the form (ax+b)^2+c.

OpenStudy (p0sitr0n):

(2x+5)^2-15

OpenStudy (anonymous):

\[ f(x)=4x^2+10x+10\\ f(x)=(2x)^2+\color{red}{2(4x)\left(\frac{10}{8}\right)}+10\\ \text{this we have to make of the form:}\\ (a+b)^2=a^2+2ab+b^2\\ \text{we obtain:}\quad b=\frac{10}{8}={5\over2}\\ f(x)=(2x)^2+2(4x)\left(5\over2\right)\color{red}{+\left(5\over2\right)^2-\left(5\over2\right)^2}+10\\ \]now, put the things together

OpenStudy (anonymous):

whoops... \[ f(x)=4x^2+10x+10\\ f(x)=(2x)^2+\color{red}{2(2x)\left(\frac{10}{4}\right)}+10\\ \text{this we have to make of the form:}\\ (a+b)^2=a^2+2ab+b^2\\ \text{we obtain:}\quad b=\frac{10}{4}={5\over2}\\ f(x)=(2x)^2+2(4x)\left(5\over2\right)\color{red}{+\left(5\over2\right)^2-\left(5\over2\right)^2}+10\\ \]

OpenStudy (anonymous):

you follow?

OpenStudy (anonymous):

Yes. Thank you. I did it the way @positron did. So my answers were all wrong.